# Fundamental Limits of Coded Polynomial Aggregation

### v1

@@ V-A Reformulation of Orthogonality
$\operatorname{rank}(\bm{U})=\min(C,K,N+1),$ (13)
as $\operatorname{rank}(\bm{U})$ == $\operatorname{rank}(\bm{V}\operatorname{diag}(\bm{w})\bm{A})$ $\leq$ $\min(\operatorname{rank}(\bm{V}$ $\operatorname{diag}(\bm{w}))$, $\operatorname{rank}(\bm{A}))$ $\overset{(a)}{=}$ $\min(\operatorname{rank}(\bm{V})$, $\operatorname{rank}(\bm{A}))$ == $\min(C$, $K$, $N+1)$, where $(a)$ follows from the fact that $\operatorname{diag}(\bm{w})$ is invertible.
@@ V-B Necessity of Theorem 2
Assume that there exist distinct data points $\{\alpha_{k}:k\in[K]\}$ satisfying the three properties in (7)–(9).
@@ V-C Sufficiency of Theorem 2
Assume $N\geq N^{*}$. We construct a set of points $\{\beta_{n}:n\in[N]\}$ satisfying the conditions in (7)-(9). The key steps of the construction are as follows. 1) We solve the linear system $\bm{U}\bm{c}=\bm{0}$. 2) We select a vector $\bm{c}\in\ker(\bm{U})$ with $c_{N}\neq 0$ and obtain the polynomial $P(z)=\frac{c_{0}}{c_{N}}+\frac{c_{1}}{c_{N}}z+\ldots+z^{N}$. From (12), the orthogonality condition in (7) is satisfied. 3) We compute the $N$ roots of the polynomial $P(z)$, i.e., $\{\beta_{n}:n\in[N]\}$.
Importantly, we will show that for a generic set of pairwise distinct $\{\alpha_{k}:k\in[K]\}$, all but a Lebesgue measure-zero subset of $\bm{c}\in\ker(\bm{U})$ lead to evaluation points $\{\beta_{n}:n\in[N]\}$ satisfying (8) and (9). For background on relevant concepts in algebraic geometry, such as affine varieties and Zariski-open dense subsets, we refer the reader to [9].
@@ V-C1 Proof of Orthogonality
Under $N\geq N^{*}$, we have $\operatorname{dim}(\operatorname{ker}(\bm{U}))>0$. Therefore, there exists a nonzero vector $\bm{c}\in\ker(\bm{U})$ such that $\bm{U}\bm{c}=\bm{0}$, as shown in equation $(c)$ of (12). Equation $(b)$ in (12) then implies that the polynomial $P_{\mathrm{form}}(z)=c_{0}+c_{1}z+c_{2}z^{2}+\ldots+c_{N}z^{N}$ satisfies the orthogonality condition $\sum_{k=0}^{K-1}w_{k}\,P_{\mathrm{form}}(\alpha_{k})\,\alpha_{k}^{j}=0$ for all $j\in[C]$. To ensure that equation $(a)$ in (12) holds, i.e., that the normalized polynomial $P(z)=\frac{P_{\mathrm{form}}(z)}{c_{N}}$ satisfies the orthogonality condition in (7), it is required that $c_{N}\neq 0$. This requirement is captured by the following lemma.
@@
The proof is sketched as follows. The set of $\bm{\alpha}=(\alpha_{0}$, $\ldots$ , $\alpha_{K-1})$ $\in$ $\mathbb{C}^{K}$ for which $\mathcal{C}_{N}$ == $\emptyset$ defines a proper affine variety in $\mathbb{C}^{K}$, and therefore has Lebesgue measure zero. Hence, for generic $\{\alpha_{k}:k\in[K]\}$, we have $\mathcal{C}_{N}\neq\emptyset$. The details are provided in Appendix A. ∎
From Lemma 2, for generic $\{\alpha_{k}:k\in[K]\}$, there exists a vector $\bm{c}\in\ker(\bm{U})$ satisfying $c_{N}\neq 0$. For such a choice of $\bm{c}$, the polynomial $P(z)=\frac{c_{0}}{c_{N}}+\frac{c_{1}}{c_{N}}z+\frac{c_{2}}{c_{N}}z^{2}+\ldots+z^{N}$ satisfies (7), and its NN roots are the desired $\{\beta_{n}:n\in[N]\}$.
@@ V-C2 Proof of Disjointness
Define a Zariski open subsets of $\ker(\bm{U})$ as $\mathcal{A}_{k}$ $\triangleq$ $\{\bm{c}\in\ker(\bm{U}):$ $\bm{A}\bm{c}]_{k} \neq$ 0}0\}, and let $\mathcal{C}_{A}$ $\triangleq$ $\bigcap_{k=0}^{K-1}$ $\mathcal{A}_{k}$. For generic $\{\alpha_{k}:k\in[K]\}$, the set $\mathcal{C}_{A}$ is nonempty.
@@
The proof is sketched as follows. Fix $k\in[K]$. The set of parameter vectors $\bm{\alpha}=(\alpha_{0},\ldots,\alpha_{K-1})\in\mathbb{C}^{K}$ for which $\mathcal{A}_{k}=\emptyset$ defines a proper affine variety in $\mathbb{C}^{K}$, and hence has Lebesgue measure zero. Therefore, for generic $\{\alpha_{k}:k\in[K]\}$, each $\mathcal{A}_{k}$ is a nonempty Zariski open subset of $\ker(\bm{U})$. Since $\ker(\bm{U})$ is a linear subspace and hence an irreducible affine variety, any finite intersection of nonempty Zariski open subsets of $\ker(\bm{U})$ is nonempty, i.e., $\mathcal{C}_{A}\neq\emptyset$. The details are provided in Appendix B. ∎
@@ V-C3 Proof of Distinction
The set of parameter vectors $\bm{\alpha}$ == $(\alpha_{0}$, $\ldots$, $\alpha_{K-1})$ $\in$ $\mathbb{C}^{K}$ for which $\mathcal{C}_{D}$ == $\emptyset$ defines a proper affine variety in $\mathbb{C}^{K}$, and therefore has Lebesgue measure zero. Hence, for generic $\{\alpha_{k}:k\in[K]\}$, we have $\mathcal{C}_{D}\neq\emptyset$. The details are provided in Appendix C. ∎
@@
Since $\ker(\bm{U})$ is an irreducible affine variety, the intersection $\mathcal{C}_{N}\cap\mathcal{C}_{A}\cap\mathcal{C}_{D}$ of the nonempty Zariski open subsets $\mathcal{C}_{N}$, $\mathcal{C}_{A}$, and $\mathcal{C}_{D}$ is nonempty and Zariski open in $\ker(\bm{U})$. Moreover, any nonempty Zariski open subset of an irreducible affine variety is dense. Hence, $\mathcal{C}_{N}\cap\mathcal{C}_{A}\cap\mathcal{C}_{D}$ is dense in $\ker(\bm{U})$. This completes the proof of Lemma 5. ∎
From Lemma 5, for generic $\{\alpha_{k}:k\in[K]\}$, all but a Lebesgue measure-zero subset of vectors $\bm{c}\in\ker(\bm{U})$ satisfy the following properties: 1) $c_{N}\neq 0$, so that the normalized polynomial $P(z)=\frac{c_{0}}{c_{N}}+\frac{c_{1}}{c_{N}}z+\ldots+z^{N}$ is well defined; 2) $\bm{A}\bm{c}]_{k}\neq 0$ for all $k\in[K]$, ensuring that $\{\alpha_{k}:k\in[K]\}$ does not contain any root of $P(z)$; and 3) $\mathrm{Disc}(P)\neq 0$, so that $P(z)$ has $N$ distinct roots. This completes the proof of sufficiency of Theorem 2.
V-D Proof of Theorem 3
Theorem 3 follows directly from the constructive proof of sufficiency in Theorem 2.

### v2

@@ V-A Reformulation of Orthogonality
$\operatorname{rank}(\bm{U})\leq\min(C,K,N+1),$ (14)
as $\operatorname{rank}(\bm{U})$ == $\operatorname{rank}(\bm{V}\operatorname{diag}(\bm{w})\bm{A})$ $\leq$ $\min(\operatorname{rank}(\bm{V}$ $\operatorname{diag}(\bm{w}))$, $\operatorname{rank}(\bm{A}))$ $\overset{(a)}{=}$ $\min(\operatorname{rank}(\bm{V})$, $\operatorname{rank}(\bm{A}))$ $\overset{(b)}{=}$ $\min(C$, $K$, $N+1)$, where $(a)$ follows from the fact that $\operatorname{diag}(\bm{w})$ is invertible. $(b)$ is due to the fact that all $\alpha_{k}$ are distinct for $k\in[K]$.
@@ V-B Necessity of Theorem 2
Assume that there exist distinct evaluation points $\{\beta_{n}:n\in[N]\}$ satisfying the three properties in (8)–(10).
@@ V-C Sufficiency of Theorem 2
Assume $N\geq N^{*}$. We construct $\{\beta_{n}:n\in[N]\}$ satisfying the conditions in (8)-(10). The key steps are as follows. 1) We solve the linear system $\bm{U}\bm{c}=\bm{0}$. 2) We select a vector $\bm{c}\in\ker(\bm{U})$ with $c_{N}\neq 0$ and obtain the polynomial $P(z)=\frac{c_{0}}{c_{N}}+\frac{c_{1}}{c_{N}}z+\ldots+z^{N}$. From (13), the orthogonality conditions in (8) is satisfied. 3) We compute the $N$ roots of the polynomial $P(z)$ as $\{\beta_{n}:n\in[N]\}$. Importantly, we will show that for a generic set of pairwise distinct $\{\alpha_{k}:k\in[K]\}$, all but a Lebesgue measure-zero subset of $\bm{c}\in\ker(\bm{U})$ lead to evaluation points $\{\beta_{n}:n\in[N]\}$ satisfying (9) and (10).
@@ V-C1 Proof of Orthogonality
Under $N\geq N^{*}$, we have $\operatorname{dim}(\operatorname{ker}(\bm{U}))>0$. Therefore, there exists a nonzero vector $\bm{c}\in\ker(\bm{U})$ such that $\bm{U}\bm{c}=\bm{0}$, as shown in equation $(c)$ of (13). Equation $(b)$ in (13) then implies that the polynomial $P_{\mathrm{form}}(z)=c_{0}+c_{1}z+\ldots+c_{N}z^{N}$ satisfies the orthogonality conditions $\sum_{k=0}^{K-1}w_{k}\,P_{\mathrm{form}}(\alpha_{k})\,\alpha_{k}^{j}=0$ for all $j\in[C]$. To ensure that equation $(a)$ in (13) holds, i.e., that the normalized polynomial $P(z)=\frac{P_{\mathrm{form}}(z)}{c_{N}}$ satisfies the orthogonality conditions in (8), it is required that $c_{N}\neq 0$. This requirement is captured by the following lemma.
@@
The proof is sketched as follows. The set of vectors $\bm{\alpha}=(\alpha_{0}$, $\ldots$ , $\alpha_{K-1})$ $\in$ $\mathbb{C}^{K}$ for which $\mathcal{C}_{N}$ == $\emptyset$ defines a proper affine variety in $\mathbb{C}^{K}$, and therefore has Lebesgue measure zero. Hence, for generic $\{\alpha_{k}:k\in[K]\}$, we have $\mathcal{C}_{N}\neq\emptyset$. See Appendix A in arXiv [16] for details. ∎
From Lemma 2, for generic $\{\alpha_{k}:k\in[K]\}$, there exists a vector $\bm{c}\in\ker(\bm{U})$ satisfying $c_{N}\neq 0$. For such a choice of $\bm{c}$, the polynomial $P(z)=\frac{c_{0}}{c_{N}}+\frac{c_{1}}{c_{N}}z+\frac{c_{2}}{c_{N}}z^{2}+\ldots+z^{N}$ satisfies (8), and its $N$ roots are the desired $\{\beta_{n}:n\in[N]\}$.
@@ V-C2 Proof of Disjointness
Define a Zariski open subset of $\ker(\bm{U})$ as $\mathcal{A}_{k}$ $\triangleq$ $\{\bm{c}\in\ker(\bm{U}):$ $\bm{A}\bm{c}]_{k} \neq$ 0}0\}, and let $\mathcal{C}_{A}$ $\triangleq$ $\bigcap_{k=0}^{K-1}$ $\mathcal{A}_{k}$. For generic $\{\alpha_{k}:k\in[K]\}$, the set $\mathcal{C}_{A}$ is nonempty.
@@
The proof is sketched as follows. The set of vectors $\bm{\alpha}$ == $(\alpha_{0}$, $\ldots$, $\alpha_{K-1})$ $\in$ $\mathbb{C}^{K}$ for which $\mathcal{A}_{k}=\emptyset$ defines a proper affine variety in $\mathbb{C}^{K}$, and hence has Lebesgue measure zero. Hence, for generic $\{\alpha_{k}:k\in[K]\}$, $\mathcal{A}_{k}$ is a nonempty Zariski open subset of $\ker(\bm{U})$. Since $\ker(\bm{U})$ is a linear subspace and hence an irreducible affine variety, any finite intersection of nonempty Zariski open subsets of $\ker(\bm{U})$ is nonempty, i.e., $\mathcal{C}_{A}\neq\emptyset$. See Appendix B in arXiv [16] for details. ∎
@@ V-C3 Proof of Distinction
The proof is sketched as follows. The set of vectors $\bm{\alpha}$ == $(\alpha_{0}$, $\ldots$, $\alpha_{K-1})$ $\in$ $\mathbb{C}^{K}$ for which $\mathcal{C}_{D}$ == $\emptyset$ defines a proper affine variety in $\mathbb{C}^{K}$, and therefore has Lebesgue measure zero. Hence, for generic $\{\alpha_{k}:k\in[K]\}$, we have $\mathcal{C}_{D}\neq\emptyset$. See Appendix C in arXiv [16] for details. ∎
@@
Since $\ker(\bm{U})$ is an irreducible affine variety, and the Zariski open subsets $\mathcal{C}_{N}$, $\mathcal{C}_{A}$, and $\mathcal{C}_{D}$ are non-empty, the intersection $\mathcal{C}_{N}\cap\mathcal{C}_{A}\cap\mathcal{C}_{D}$ is nonempty and Zariski open in $\ker(\bm{U})$. Moreover, any nonempty Zariski open subset of an irreducible affine variety is dense. Hence, $\mathcal{C}_{N}\cap\mathcal{C}_{A}\cap\mathcal{C}_{D}$ is dense in $\ker(\bm{U})$. This completes the proof of Lemma 5. ∎
From Lemma 5, for generic $\{\alpha_{k}:k\in[K]\}$, all but a Lebesgue measure-zero subset of vectors $\bm{c}\in\ker(\bm{U})$ satisfy the following properties: 1) $c_{N}\neq 0$, so that the normalized polynomial $P(z)=\frac{c_{0}}{c_{N}}+\frac{c_{1}}{c_{N}}z+\ldots+z^{N}$ is obtained and satisfies orthogonality conditions; 2) $\bm{A}\bm{c}]_{k}\neq 0$ for all $k\in[K]$, ensuring that $\{\alpha_{k}:k\in[K]\}$ does not contain any root of $P(z)$; and 3) that the $N$ roots of $P(z)$ are distinct.
This completes the proof of the sufficiency of Theorem 2, and hence completes the proof of Theorem 2.